2.)What is investment in a closed economy if you have the following economic data?
Y = $10 trillion
C = $5 trillion
TR = $2 trillion
G = $2 trillion

  1. A.

    $2 trillion

    B.

    $3 trillion

    C.

    $5 trillion

    D.

    cannot be determined without information on taxes (T)

4 points   
Answer #1

B) $3 trillion

Y = GDP, C = Consumption, G = Government purchases, T = Taxes, and TR = Transfers

Y = C + I + G ( I = Investment )

I = Y - C - G = 10 - 5 - 2 = $3 trillion

public saving plus private saving is equal to Investment